An oil company estimates that only 1 well in 15 will yield commercial quantities of oil. Assume that successful drilled wells represent independent events. If 11 wells are drilled, find the probability of obtaining a commercially successful well for the following number of times: exactly 2

(Round your answer to six decimal places.)
Answer #1

The probability distribution of number of wells yielding commercial quantities of oil follows binomial distribution.

P(a well will yield commercial quantities of oil), p = 1/15

P(a well will not yield commercial quantities of oil), q = 1 - 1/15 = 14/15

Number of wells drilled, n = 11

P(X) = nCx px qn-x

P(exactly 2 well out of 15 yielding commercially successful quantities of oil) = P(2)

= 11C2 x (1/15)2 x (14/15)9

= 0.131375